Webbie10000:

Yuri bakes lemon bars in a pan shaped like a right rectangular prism. The volume of the pan is 150 in³. The width of the pan is 7 and 1/2 in., and its height is 2 in. What is the length of the pan? Enter your answer in the box. in.

Vocaloid:

volume = width * length * height you are given volume, width, and height, plug in and solve for length

Webbie10000:

answer pls

Vocaloid:

We have a rule against direct answers. Please think about what I have said and how it applies to your problem.

Webbie10000:

ummm i can't do math i struggle alot

AnimeGhoul8863:

Explanation: _____________________________________________ Volume of a rectangular prism:   V = L * w * h ;  in which:  V = volume = 150 in.³ ;  L = length = ? ; w = width = 7.5 in. ;  h = height = 2 in.  _______________________________________________ Given these values, we want to solve for "L" (the "length") ;  _______________________________________________ Rewrite the formula/equation:  _______________________________________________  V = L * w * h ; by isolating "L" (the "length"); on one side of the equation;   by dividing BOTH sides of the equation by "(w * h)" ; _______________________________________________ V / (w * h) = (L * w * h) / (w *h) ; to get: ________________________________________________    V / (w * h) = L ;  ↔  L = V / (w * h) ; ________________________________________________ Plug in our given values; and solve for "L" ;  ________________________________________________        L = 150 in.³ / (7.5 in. * 2 in.) ;     →  L = 150 in.³ / (15 in.²)  ;      →  L = (150 / 15) in. ;        →  L = 10 in. ; ______________________________________________

AnimeGhoul8863:

so whats the length of the pan

Vocaloid:

If something is unclear you are welcome to ask questions or clarify ^^ anyway, we know that volume = width * length * height you can replace "volume" with 150, height with 2, and width with (7 1/2) so 150 = (2)(7 1/2)x solve for x

Webbie10000:

what i dont understand anime

Webbie10000: